# Community Discovery Algorithm Based on Spatio-temporal Graph Embedding in Dynamic Social Networks

### v1

@@ 4.3.3 Comparative Analysis of Community Discovery Methods
In Fig. 5 we present the evaluation results of community discovery algorithms on dynamic networks.
Figure 5: The average evaluation results of all algorithms in datasets: the average purity in synthetic datasets (a), the average NMI in synthetic datasets (b), the average purity in real datasets (c), and the average NMI in real datasets (d).
From the figure, it is clear that CDA-SGE shows a noticeable effect on community partitioning for both real and synthetic network datasets. However, its performance on large real dynamic networks is significantly lower than on smaller ones, highlighting that the complexity and variability of large dynamic networks remain key challenges in the research. However, CDA-SGE has achieved strong evaluation results across both artificially synthesized and real dynamic network datasets. The accuracy of purity and NMI in various datasets are presented in Tab. 4.
@@
DBLP5 48.10 37.60
Table 4: Accuracy of Purity and NMI in Various Datasets
@@
Comparing the proposed algorithm with traditional static community discovery algorithms, such as Louvain, and dynamic community discovery algorithms such as Dyperm and FacetNet, CDA-SGE outperforms these methods in community partitioning. FacetNet’s performance on artificial networks is suboptimal, although it performs better on real networks, suggesting that it focuses more on changes in real-world social activities. Dyperm demonstrates relatively stable community partitioning performance, while Louvain performs well in real online communities.
Figure 6: Number of communities divided by the thesis algorithms in PS dataset.
@@
The comparison between CDA-SGE and CDA-SGE-K primarily evaluates the impact of clustering algorithms on the partitioning results of the graph embedding model. From the results, it is observed that the evaluation results for both the SOM and the K-means clustering algorithms are similar, with differences within 5%. The SOM algorithm was chosen because it does not require specifying the number of clusters in advance, only the number of competing units, neighborhood range, and learning rate. This allows for adaptive community division based on the embedding representation of nodes. In contrast, the K-means algorithm requires prior knowledge of the number of clusters, which is often difficult to determine in real dynamic networks. Hence, SOM clustering is more suitable for this work.
Figure 7: The evaluation results of the thesis algorithms in PS dataset: the average purity in PS (a), and the average NMI in PS (b).
@@
In Fig. 6, we present the comparison of community partitioning results for the PS dataset, while in Fig. 7, we present the evaluation results of both algorithms on the same dataset. From Fig. 6, it is evident that CDA-SGE’s partitioning results are closely aligned with the actual number of communities, with only minor discrepancies. In contrast, DyCTWE shows a maximum difference of 5 communities from the real partitioning. Fig. 7 reveals that the quality of DyCTWE’s partitioning deteriorates over time, indicating error accumulation and a decline in community discovery performance. This issue likely arises because DyCTWE overly emphasizes local incremental updates, neglecting the overall evolution of the network. On the other hand, CDA-SGE demonstrates an increasing trend in performance over time, suggesting that spatiotemporal graph embedding effectively captures the dynamic evolution of nodes and offers robust representation capabilities for dynamic networks.
@@ 5 Conclusion
In this work, we explore the complex structure and rules of social networks in the spatiotemporal dimension through a community discovery algorithm based on spatiotemporal graph embedding. Experimental validation demonstrates the effectiveness of the proposed algorithm in mining dynamic social network communities, offering new insights and methods to understand group organization and information flow in such networks.
For future work, we plan to optimize the algorithm by integrating advanced deep learning architectures, such as attention-based neural networks, to improve spatiotemporal feature learning, automatically adjust embedding parameters, and enhance the accuracy of community discovery. Additionally, we plan to extend the algorithm to dynamic network analysis in more fields, such as transportation networks, biological networks, particle interactions, etc., to verify its generality and effectiveness. In addition, we will integrate the CDA-SGE with other community discovery algorithms, complement each other’s strengths and weaknesses, and build a more powerful hybrid algorithm framework to cope with the increasingly complex and changing network environment, providing stronger technical support for the development of dynamic social network research and related application fields.

### v2

@@ 4.5.3 Comparative Analysis of Community Discovery Methods
In Fig. 5(a) and  5(d) we present the evaluation results of community discovery algorithms on dynamic networks.
(a) Average purity on Synthetic Datasets
(b) Average NMI on Synthetic Datasets
(c) Average Purity on Real Datasets
(d) Average NMI on Real Datasets
Figure 5: Comprehensive comparison of community discovery algorithms. GTENN demonstrates superior performance over traditional static (Louvain), dynamic (Dyperm, FacetNet), and embedding-based (DyCTWE) methods across both synthetic and real datasets.
From these figures, it is clear that GTENN shows a noticeable effect on community partitioning for both real and synthetic network datasets. However, its performance on large real dynamic networks is significantly lower than on smaller ones, highlighting that the complexity and variability of large dynamic networks remain key challenges in the research. However, GTENN has achieved strong evaluation results across both artificially synthesized and real dynamic network datasets. The evaluation metrics, purity and NMI in various datasets, are presented in Tab. 5.
@@
DBLP 48.11 37.60
Table 5: Purity and NMI in the synthetic and real datasets of the proposed GTENN alogrithm
@@
Comparing the proposed algorithm with traditional static community discovery algorithms, such as Louvain, and dynamic community discovery algorithms such as Dyperm and FacetNet, GTENN outperforms these methods in community partitioning. FacetNet’s performance on artificial networks is suboptimal, although it performs better on real networks, suggesting that it focuses more on changes in real-world social activities. Dyperm demonstrates relatively stable community partitioning performance, while Louvain performs well in real online communities.
Figure 6: Number of communities identified by GTENN and DyCTWE on the PS dataset, compared with the ground-truth community structure across time steps.
@@
The comparison between GTENN and GTENN-K primarily evaluates the impact of clustering algorithms on the partitioning results of the graph embedding model. From the results, it is observed that the evaluation results for both the SOM and the K-means clustering algorithms are similar, with differences within 5%. The SOM algorithm was chosen because it does not require specifying the number of clusters in advance, only the number of competing units, neighborhood range, and learning rate. This allows for adaptive community division based on the embedding representation of nodes. In contrast, the K-means algorithm requires prior knowledge of the number of clusters, which is often difficult to determine in real dynamic networks. Hence, SOM clustering is more suitable for this work.
(a) Average purity in PS
(b) Average NMI in PS
Figure 7: Evaluation results of GTENN and DyCTWE on the PS dataset over time. The plots report Purity (left) and NMI (right) at each time step, showing that GTENN maintains consistently higher performance across the dynamic snapshots.
@@
In Fig. 6, we present the comparison of community partitioning results for the PS dataset, while in Fig. 7, we present the evaluation results of both algorithms on the same dataset. From Fig. 6, it is evident that GTENN’s partitioning results are closely aligned with the actual number of communities, with only minor discrepancies. In contrast, DyCTWE shows a maximum difference of 5 communities from the real partitioning. Fig. 7 reveals that the quality of DyCTWE’s partitioning deteriorates over time, indicating error accumulation and a decline in community discovery performance. This issue likely arises because DyCTWE overly emphasizes local incremental updates, neglecting the overall evolution of the network. On the other hand, GTENN demonstrates an increasing trend in performance over time, suggesting that spatiotemporal graph embedding effectively captures the dynamic evolution of nodes and offers robust representation capabilities for dynamic networks.
4.6 Comparison with Attention-Based Methods
In this section, we present the evaluation results in terms of purity and NMI and computing power consumption agains some attention-based algorithms, including ASTGCN [28], MSTGCN [29] and ST-GAT [30].
The results in Table 6 demonstrate that the proposed GTENN algorithm exhibits outstanding performance in both community detection quality and computational efficiency across two representative datasets: LFR2 and PS.
Table 6: Comprehensive Performance Comparison of Different Algorithms
Algorithm Dataset Purity NMI Training Time Memory/Params
GTENN LFR2 0.937 0.949 12.3 2.1GB/48K
PS 0.927 0.901 8.5 1.8GB/48K
NWFCM LFR2 0.721 0.698 6.2 0.8GB/12K
PS 0.704 0.681 4.1 0.6GB/12K
ASTGCN LFR2 0.834 0.821 28.6 4.5GB/156K
PS 0.798 0.785 18.7 3.8GB/156K
MSTGCN LFR2 0.856 0.843 45.3 6.8GB/243K
PS 0.841 0.827 32.1 5.9GB/243K
ST-GAT LFR2 0.812 0.795 22.4 3.7GB/128K
PS 0.789 0.773 16.2 3.1GB/128K
The proposed GTENN achieves Purity values of 0.937 and 0.927 on the LFR2 and PS datasets, respectively, with corresponding NMI values of 0.949 and 0.901, significantly outperforming all baseline algorithms. Compared to the second-best performing MSTGCN, GTENN shows an average improvement of approximately 8.2% in Purity and 9.1% in NMI. This indicates that GTENN, through its spatiotemporal joint modeling strategy combining GCN and GRU, can more accurately capture community evolution patterns in dynamic networks than attention-based approaches.
The performance comparison reveals distinct characteristics of different algorithm categories:
  • Traditional methods like NWFCM [31] show the lowest performance but maintain computational efficiency.
  • Attention-based methods (ASTGCN, MSTGCN, ST-GAT) achieve moderate performance but with significant computational overhead.
  • Our spatiotemporal approach achieves the best performance while maintaining reasonable computational cost.
GTENN also demonstrates clear efficiency advantages over attention-based methods. Although its training time is slightly higher than that of the traditional NWFCM approach, it remains substantially lower than that of more complex attention-based models. Specifically, compared with MSTGCN, GTENN reduces training time by approximately 70
The robustness of GTENN is further demonstrated by its stable performance across datasets of different nature and complexity. From the synthetic LFR2 to the real-world PS network, the decline in Purity is less than 2%, while attention-based algorithms typically experience 3-8% decreases. This demonstrates the superior generalization capability and stability of the proposed algorithm. The consistent performance across different dataset characteristics (synthetic vs. real-world, different scales and domains) indicates that our spatiotemporal modeling approach is more robust than attention mechanisms that may be sensitive to specific network structures.
The results further underscore the advantages of GTENN over attention-based methods. Specifically, ASTGCN, although effective for traffic prediction, shows limited suitability for community detection tasks. MSTGCN leverages multi-view attention mechanisms but incurs substantial computational overhead, making it less practical for large or complex networks. ST-GAT integrates graph attention networks for spatiotemporal modeling, yet its performance exhibits instability across different datasets. In conclusion, the proposed GTENN achieves superior detection accuracy measured in purity and NMI, and computational efficiency, offering more stable and effective spatiotemporal modeling than attention-based counterparts, while maintaining substantially lower computational complexity.
@@ 5 Conclusion
In this work, we proposed GTENN, a community discovery algorithm based on spatiotemporal graph embedding, to explore the complex structure and dynamics of social networks. Experimental validation demonstrates that GTENN effectively uncovers community evolution in dynamic networks, providing new insights into group organization and information flow.
A key characteristic of GTENN is its discrete-time snapshot framework, which is well suited to datasets collected in aggregated intervals. While continuous-time approaches such as TGN and JODIE offer fine-grained event-level modeling, our datasets are not inherently event-driven, and converting them into event streams could distort their temporal structure. For this reason, we adopt the snapshot-based representation, while extending GTENN to continuous-time settings with appropriate datasets remains an important direction for future research.
Looking ahead, we plan to enhance GTENN with advanced deep learning architectures, such as attention-based networks, to strengthen spatiotemporal feature learning and improve performance across evaluation metrics. We will also extend the framework to diverse domains, including transportation, biology, and particle interactions, to test its generality. Of particular significance is the financial sector, where applications in fraud detection and anti–money laundering (AML) can uncover hidden transactional patterns and anomalous group behaviors that traditional methods often miss. Such extensions would not only advance domain-specific impact but also demonstrate the broader practical utility of GTENN for understanding complex and evolving networks.
Acknowledgment
The authors gratefully acknowledge the support of Qilu Bank, which provided the necessary CPU and GPU machine time for the computational experiments. We also wish to express our gratitude to its Financial Mathematical Lab for providing a conducive research environment and valuable resources.
Data Availability Statement
This study utilized publicly available open datasets, including real-world dynamic network data and benchmark datasets. The synthetic datasets generated and analyzed during the study are available from the corresponding author upon reasonable request.